OpenStudy (anonymous):

Help-I need to graph sin and cos type 2 y=1/2sin(1/2x) 0 | 0 pi/2 | 2 pi | 0 3pi/2| -2 2pi | 0 For some reason I am wrong? Can someone help me.

OpenStudy (dumbcow):

amplitude = 1/2 not 2 and period is 4pi not 2pi

OpenStudy (dumbcow):

period = 2pi/(1/2) = 4pi

OpenStudy (anonymous):

so how would it be graphed? this is what my graph looks like.

OpenStudy (dumbcow):

\[y = A \sin(Bx)\] Period = 2pi/B 0 | 0 P/4 | A P/2 | 0 3P/4 | -A P | 0

OpenStudy (anonymous):

To draw the graph, plot all x -intercepts, minima, and maxima within one period. It says I'm wrong. I've literally been stuck on these sorts of problems for days.

OpenStudy (dumbcow):

your amplitude is right, but your period is still off B = 1/2 Period = 2pi/B = 2pi/(1/2) = 4pi

OpenStudy (dumbcow):

0 | 0 Period/4 | A Period/2 | 0 3/4 Period | -A Period | 0